OpenStudy (anonymous):

An infinite geometric series with second term 9/4 has a sum of 81/8. Determine the possible values for the first term. I understand the formula for an infinite geometric series is \[S=\frac{a}{1-r}\] Given that, I need to find a and r. I thought I could find r by using \[t_n=ar^{n-1}\] but I still do not know r or a. I would know how to do this if I was given two terms, but not when I am given only one.

OpenStudy (anonymous):

\(t_n=ar^{n-1} \) for \( n =2, \frac 94 = ar \) and \[ S=\frac{a}{1-r} \implies 81(1-r) = 8a \implies r =1- \frac 8 {81}a\]

OpenStudy (anonymous):

Substituting, \[\frac 9 4 = a - \frac 8 {81}a^2 \implies \frac 8 {81} a^2-a-\frac94 =0 \]

OpenStudy (anonymous):

And all of these assuming my algebra is right :)

OpenStudy (anonymous):

Thank you!! :)

OpenStudy (anonymous):

Glad to help :)